### v3

# Ziliang Zhang College of Oceanic and Atmospheric Sciences

Abstract
Traditional wave forecasting models, although based on energy conservation equations, are computationally expensive. On the other hand, existing deep learning geophysical fluid models, while computationally efficient, often suffer from issues such as energy dissipation in long-term forecasts. This paper proposes a novel energy-balanced deep learning wave forecasting model called OceanCastNet (OCN). By incorporating wind fields at the current, previous, and future time steps, as well as wave fields at the current and previous time steps as input variables, OCN maintains energy balance within the model. Furthermore, the model employs adaptive Fourier operators as its core components and designs a masked loss function to better handle the impact of land-sea boundaries. A series of experiments on the ERA5 dataset demonstrate that OCN can achieve short-term forecast accuracy comparable to traditional models while exhibiting an understanding of the wave generation process. In comparative experiments under both normal and extreme conditions, OCN consistently outperforms the widely used WaveWatch III model in the industry. Even after long-term forecasting, OCN maintains a stable and energy-rich state. By further constructing a simple meteorological model, OCN-wind, which considers energy balance, this paper confirms the importance of energy constraints for improving the long-term forecast performance of deep learning meteorological models. This finding provides new ideas for future research on deep learning geophysical fluid models.
Keywords Ocean Wave Forecasting  $\cdot$ Deep Learning  $\cdot$ Numerical Modelling
1 Introduction
Traditional geophysical forecasting models are mostly based on conservation principles [1, 2]. In atmospheric and oceanic models, this conservation is reflected in the conservation of momentum and mass, while in ocean wave models, it is represented by the conservation of energy [3]. Traditional models discretize the equations in these conservation laws and perform a large number of conserved quantity calculations in each minimal region of the ocean or atmosphere, thereby simulating and forecasting the overall process. In pursuit of accurate process simulations, these models continuously refine the computational domains, resulting in enormous computational requirements, often necessitating extensive computational resources and time. Consequently, only forecasting agencies equipped with high-performance computers can conduct the corresponding forecasts.
In recent years, deep learning meteorological models have achieved fruitful results. Since the proposal of FourCastNet, deep learning meteorological models such as Pangu-Weather, GraphCast, and FUXI have demonstrated not only computational speeds far surpassing traditional models but also forecast accuracies that have improved from being comparable [4] to comprehensively surpassing traditional models [5, 6]. In the oceanic domain, models like AI-GOMS and XIHE have begun to apply deep learning to ocean forecasting. Among them, AI-GOMS has also attempted wave forecasting and achieved promising results [7]. However, AI-GOMS has a temporal resolution of 1 day, which cannot provide ideal guidance for marine operators, thus limiting its practical application. Moreover, the forecast duration of current deep learning oceanic models is generally short, and their understanding and simulation capabilities for long-term processes need further improvement.
It is noteworthy that the aforementioned deep learning models are primarily constructed based on statistical methods such as autoregression rather than conservation physical laws like traditional models. Due to the lack of energy input, this leads to issues in long-term forecasting, with the most obvious being the energy dissipation problem in the forecast results. For example, after long-term forecasting, the results of FourCastNet often exhibit significant energy decay and even numerical collapse.
This paper first constructs an energy-balanced deep learning wave model. This construction approach enables the model to achieve short-term forecast accuracy comparable to traditional models while also demonstrating the model’s understanding of the wave generation process. Even when the initial wave field is zero, the model can generate a wave field with a zonal mean anomaly correlation coefficient (ACC) higher than 0.8 solely based on wind field input and maintain stable ACC and root mean square error (RMSE). After long-term forecasting, the model still maintains a state of sufficient energy. To further validate our idea, we performed a simple energy balance modification on FourCastNet and compared the long-term forecast results. We observed that when FourCastNet’s long-term forecast results exhibited energy dissipation or even numerical collapse, the modified model maintained a stable and energy-rich state. This confirms that our idea is correct.
2 Data and Methods
The wave process is mainly composed of wind sea and swell. Wind sea is directly driven by the interaction between wind and the sea surface in the wind area, while swell is formed when wind sea leaves the wind area [8]. To capture the spatial-temporal evolution patterns of waves while satisfying the energy conservation requirement, we select the wind and wave fields at the current and previous time steps, together with the wind field at the future time step, as the model’s input variables. Specifically, the input-output relationship of the model can be represented by the following function:
$h(t+1)=f[u_{10}(t-1),v_{10}(t-1),h(t-1),u_{10}(t),v_{10}(t),h(t),u_{10}(t+1),v_{10}(t+1)],$ (1)
where $\mathit{h}(t)$ represents the significant wave height at time $t$, $\mathit{u}{10}(t)$ and $\mathit{v}{10}(t)$are the meridional and zonal wind speeds at 10 m above the sea surface, respectively, and $f$ represents the model’s mapping function. We name the model OceanCastNet (OCN).
In traditional wave models, the governing equations usually consist of multiple partial differential equations. Considering the excellent performance of Fourier neural operators in solving partial differential equations [9], such as the ability to directly parameterize integration kernels in Fourier space and achieve efficient solutions to partial differential equations, we adopt Adaptive Fourier Neural Operators (AFNO) as the core components of the model [10]. Compared to atmospheric models, the presence of land has a significant impact on the training of wave models, such as causing the model to overly focus on land areas or the occurrence of land errors canceling out ocean errors. To make the model more focused on ocean area forecasting, we design a masked loss function. Specifically, we introduce a mask matrix corresponding to the land-sea distribution, so that the errors in land areas are masked out during backpropagation, making the model more focused on ocean area forecasting.
In model training, we use the ERA5 reanalysis data from the European Centre for Medium-Range Weather Forecasts (ECMWF). ERA5 provides global hourly 10-m wind field data at 0.25-degree resolution and hourly significant wave height data at 0.5-degree resolution. These data integrate multi-source observation data such as conventional observations, satellite scatterometers, and satellite altimeters, and are thus very close to the real situation [11]. To align with existing deep learning weather models and avoid the impact of sea ice changes on the size of ocean areas, we use a 6-hour temporal resolution and select the latitude range from -61° to 61°. We use 41 years of ERA5 data from 1980 to 2020, with 1980-2017 as the training set, 2018-2019 as the validation set, and 2020 as the test set. Model training is conducted on 32 NVIDIA Tesla V100 GPUs, with a training time of 16 hours.
To evaluate the model’s forecasting performance, we use the forecasting results of the University of Hawaii’s WaveWatch III (WW3) global wave model as a benchmark for comparison. WW3 is driven by the wind field data from NOAA/NCEP’s Global Forecast System (GFS) and provides hourly global wave forecasting data at 0.5-degree resolution. Both WW3 and ERA5 are widely used high-quality wave datasets in the industry and can provide reliable references for model performance evaluation.
We adopt ACC and RMSE as the main evaluation metrics. ACC measures the correlation between the model’s forecasted anomalies and the actual anomalies, while RMSE measures the overall bias of the model’s forecasts. By taking the zonal mean, we can eliminate the spatial variability of the wave field, making the evaluation metrics more focused on the model’s ability to depict the global mean state; and by anomaly processing, we can filter out the seasonal cycle signal in the wave field, making the evaluation metrics more focused on the model’s forecasting level of wave anomalies. The detailed formulas for calculating ACC and RMSE are provided in the Appendix A.
3 Experiments and Results
Our experiments consist of three parts. In the first part, we validate the accuracy of the model’s forecasts by using ERA5 data as the forcing field (i.e., the input conditions for the model). In the second part, we use the wind field data predicted by the model as the forcing field and compare it with traditional models to verify the improvement in the model’s forecast performance. In the third part, we use unconventional wave fields as input to verify the physical characteristics of the model. Each experiment is divided into two scenarios: extreme and normal. The extreme scenario evaluates the forecast results in the event area during the occurrence of the strongest tropical cyclone (typhoon or hurricane) of the year. The normal scenario randomly selects multiple events (e.g., 20 events) for forecasting and calculates the average results.
3.1 Performance Evaluation on ERA5 Dataset
We employ ACC and RMSE as evaluation metrics, where ACC measures the correlation between forecasts and actual values, and RMSE quantifies the magnitude of forecast errors. The evaluation is divided into two parts: model performance assessment under normal conditions and extreme situations. To verify the model’s forecasting performance in extreme situations, we select the main impact periods of Super Typhoon Goni and Hurricane Iota, the strongest tropical cyclones in the Pacific and Atlantic Oceans, respectively, in 2020. For model performance assessment under normal conditions, we conduct forecasts for the 3rd to the 18th day of each month and take the average of these 16-day forecast results.
3.1.1 normal conditions
We simulate a 15-day period from the 3rd to the 18th of each month in 2020. The simulation process uses the ERA5 wind field and wave field at 12:00 and 16:00 on the 2nd of each month as the initial time step, with the ERA5 wind field as the forcing field for continuous forecasting. The forecast results are then subtracted from the actual values to obtain the error distribution. Figure 1 shows the averaged forecast results.
Figure 1: a) Error distribution of the OceanCastNet (OCN) model at the 1st time step in predicting wave heights. The errors are visualized on a world map, with red indicating overestimation and blue indicating underestimation. c) Error distribution of the OCN model at the 59th time step in predicting wave heights. The error patterns are more pronounced compared to the 1st time step, suggesting an increase in prediction uncertainty over time. b) Accuracy (ACC) curve of the OCN model as a function of the prediction time step. The ACC measures the overall agreement between the predicted and true wave heights. The model exhibits high accuracy in the early time steps, which gradually decreases as the prediction horizon extends. d) Root Mean Square Error (RMSE) curve of the OCN model as a function of the prediction time step. The RMSE quantifies the average magnitude of the prediction errors. The model achieves low RMSE values in the initial time steps, indicating good performance, but the RMSE increases over time due to the accumulation of errors. The results demonstrate that the OCN model maintains high accuracy and low errors in the early stages of prediction, but its performance progressively declines as the prediction time step increases, reflecting the growing uncertainty in long-term wave forecasting.
During the 59 time steps (14.75 days) of the forecasting process, the model’s mean absolute error(MAE) is 0.08 m, with a minimum ACC of 0.82 and a maximum RMSE of 0.45 m. Considering the evaluation curves, the model performs well in forecasting the wave field. It is worth noting that although the model’s error decreases over time, it does not continue to decline like other forecasting models. Instead, it stabilizes at around 0.82 after the initial decrease. Based on this, we hypothesize that the model has learned the physical properties of waves.We will present relevant experiments in Section 3.3.
3.1.2 extreme situations
The strongest typhoon in 2020 was Typhoon Goni, which formed on October 26, 2020, and was upgraded to a typhoon on October 29 after a period of rapid intensification. The typhoon eventually dissipated on November 6. Therefore, we selected the ERA5 wave and wind fields at 12:00 and 16:00 on October 28 as the initial conditions, and the ERA5 wind field as the forcing field for continuous forecasting. The forecasting results and the corresponding metric curves are shown in Figure 2.
Figure 2: Performance of the OCN model during the generation and dissipation of Typhoon Goni. a) and d) show the 10-meter wind speed at the sea surface at time steps 1 and 13, corresponding to the formation and the strongest stage of Typhoon Goni, respectively. b) and e) display the error distribution of the OCN model at these two time steps. During the formation stage of Typhoon Goni, the model has larger errors near the typhoon center, while at its strongest stage, the error distribution is relatively uniform, and the overall error has decreased. c) and f) present the ACC and RMSE curves of the OCN model throughout the simulation period. Similar to Hurricane Iota, the ACC generally shows an upward trend, and the RMSE exhibits a downward trend, but the magnitude of change is relatively smaller, indicating that the improvement in the prediction performance of the OCN model is relatively gradual during the lifecycle of Typhoon Goni.
During the 35-time-step (8.75 days) forecasting period, within the project area (longitude 120°E to 170°E, latitude 0°N to 50°N), the MAE was 0.24 m, the minimum ACC was 0.98, and the maximum RMSE was 0.32 m. Considering the evaluation curves, the model’s forecasting performance during Typhoon Goni can be regarded as satisfactory.
The strongest hurricane in the Atlantic region in 2020 was Hurricane Iota, which formed on November 13, 2020, and dissipated on November 18. Therefore, we selected the ERA5 wave and wind fields at 12:00 and 16:00 on November 12 as the initial conditions, and the ERA5 wind field as the forcing field for continuous forecasting. The forecasting results and the corresponding metric curves are shown in Figure 3.
Figure 3: Performance of the OCN model during the generation and dissipation of Hurricane Iota. a) and d) show the 10-meter wind speed at the sea surface at time steps 1 and 15, corresponding to the formation and the strongest stage of Hurricane Iota, respectively. b) and e) display the error distribution of the OCN model at these two time steps, revealing larger errors near the hurricane center, but the overall error decreases over time. c) and f) present the ACC and RMSE curves of the OCN model throughout the simulation period. The OCN curve generally shows an upward trend, while the RMSE curve exhibits a downward trend, indicating that the prediction performance of the OCN model improves continuously during the lifecycle of Hurricane Iota.
During the 23-time-step (5.75 days) forecasting period, within the forecasting area (longitude 100°W to 60°W, latitude 0°N to 20°N), the MAE was 0.21 m, the minimum ACC was 0.84, and the maximum RMSE was 0.33 m. Considering the evaluation curves, the model’s forecasting performance in the hurricane region is also highly satisfactory.
3.2 Performance Comparison against Traditional Forecasting Models
In wave forecasting, WW3 model is one of the most powerful traditional forecasting models. As mentioned earlier, WW3 is a third-generation wave model based on the energy balance equation. For the comparison data, we used WW3 Global Wave Model data from the University of Hawaii, which is driven by NOAA/NCEP’s Global Forecast System (GFS) winds. This dataset provides hourly wave forecasting data at 0.5-degree spatial resolution. When comparing with the OCN model, our model’s initial fields were derived from ERA5 data, and the wind fields used for forcing were the same as those used in the WW3 model. As model errors accumulate over time, shorter forecast lead times generally result in higher model accuracy. The University of Hawaii’s WW3 model updates its forecasts once a day and only retains the forecast data for the next 24 hours, as this period has the smallest forecast error and the highest reliability. Therefore, when comparing with the OCN model, we selected this ’24-hour short-term forecast’ result from the WW3 model. Since the University of Hawaii’s model started forecasting in 2023, we selected the main impact periods of the strongest typhoon (Mawar) and the strongest hurricane (Lee) in 2023 to validate the model’s forecasting performance. Additionally, we randomly selected one day from each month, excluding these two periods, and compared the model’s performance based on the average results. As the ERA5 reanalysis data for December 2023 was not fully updated at the time of writing this article, we only sampled from the first 11 months.
3.2.1 normal conditions
We selected the following dates as the starting points for forecasting: January 27, February 10, March 18, April 12, May 6, June 24, July 13, August 6, September 26, October 2, and November 3, 2023. For each starting point, we conducted a 42-hour (7 time steps) forecast. Subsequently, we calculated the absolute difference between the forecasted results and the ERA5 data, and then averaged the differences across all the selected dates. The forecasting results are shown in Figure 4.
Figure 4: Comparison of absolute errors between the OCN and WaveWatch III (WW3) models for the same event. a) and d) show the global absolute error distribution of the OCN model at time steps 1 and 8 (the last step), respectively. b) and e) display the global absolute error distribution of the WW3 model at the same time steps. By comparison, it can be observed that the errors of OCN are generally smaller than those of WW3, especially in the Southern Hemisphere and the North Atlantic region. c) presents the comparison curve of the ACC between the two models over the 8 time steps. The ACC of OCN is consistently higher than that of WW3, indicating that its predictions have a better correlation with the true values. f) shows the comparison curve of the RMSE between the two models. The RMSE of OCN is consistently lower than that of WW3, suggesting that its prediction errors are smaller. It is worth noting that since the results of WW3 are updated every 24 hours, only 8 time steps are selected for comparison here.
During the first 7 time steps (1.75 days) of the forecasting process, the MAE of the OCN model was 0.14, which was lower than the MAE of the WW3 model (0.30). In terms of ACC, the minimum value for the OCN model was 0.99, higher than the 0.92 of the WW3 model. Regarding RMSE, the maximum value for the OCN model was 0.36, lower than the 1.26 of the WW3 model. Considering the evaluation curves, we can conclude that the short-term forecasting performance of the OCN model under normal conditions is better than that of the WW3 model.
3.2.2 extreme situations
Typhoon Mawar (2023) was the strongest typhoon in the Pacific region in 2023, forming on May 19 and dissipating on June 11. It reached its peak intensity of 68 m/s on May 26. Therefore, we selected May 25 as the starting point for forecasting, using the ERA5 wave and wind fields at 12:00 and 18:00 on May 24 as the initial conditions, and the GFS data as the wind forcing. We conducted a 3-day forecast and compared the differences between the OCN model and the WW3 model. Some of the results are shown in Figure 5.
Figure 5: Comparison of the performance of OCN and WW3 models during Typhoon Mawar. As the WW3 results are updated every 24 hours, this figure only selects 11 time steps covering the strongest period (3 days) of the typhoon for comparison. a) and e) show the 10-meter wind speed at the sea surface at time steps 1 and 11, corresponding to the formation and the strongest stage of Typhoon Mawar, respectively. b) and f) display the error distribution of the OCN model at these two time steps, while c) and g) show the error distribution of the WW3 model at the same time steps. By comparison, it can be observed that the errors of OCN are generally smaller than those of WW3, especially near the typhoon center. d) presents the ACC between the two models over the 11 time steps. The ACC of OCN is consistently higher than that of WW3, indicating that its predictions have a better correlation with the true values. (h) shows the comparison curve of the RMSE between the two models. The RMSE of OCN is consistently lower than that of WW3, suggesting that its prediction errors are smaller. Overall, OCN demonstrates higher prediction accuracy than WW3 in the simulation of Typhoon Mawar.
During the 11-time-step (2.75 days) forecasting period, within the forecasting area (longitude 120°E to 162°E, latitude 0°N to 30°N), the MAE of the OCN model (0.16 m) was lower than that of the WW3 model (0.26 m). In terms of ACC, the minimum value for the OCN model (0.98) was higher than that of the WW3 model (0.97). Regarding RMSE, the maximum value for the OCN model (0.51 m) was lower than that of the WW3 model (0.71 m). Considering the evaluation curves, we can conclude that the OCN model performed better than the WW3 model in forecasting the Mawar region.
Hurricane Lee (2023) was the strongest hurricane in the Atlantic region in 2023, forming on September 5 and dissipating on September 18. It rapidly intensified into a Category 5 hurricane on September 7, with its wind speed increasing by 85 mph (140 km/h) within 24 hours, becoming the third-fastest rapid intensification on record in the Atlantic. Rapid intensification refers to a significant increase in a tropical cyclone’s maximum sustained winds within a short period, typically 24 hours or less. We selected September 6 as the starting point for forecasting, using the ERA5 wave and wind fields at 12:00 and 18:00 on September 5 as the initial conditions, and the GFS data as the wind forcing. We conducted a 3-day forecast and compared the differences between the OCN model and the WW3 model. Some of the results are shown in Figure 6.
Figure 6: Comparison of the performance of OCN and WW3 models during Hurricane Lee. As the WW3 results are updated every 24 hours, this figure only selects 11 time steps covering the strongest period (3 days) of the hurricane for comparison. a) and e) show the 10-meter wind speed at the sea surface at time steps 1 and 11, corresponding to the formation and the strongest stage of Hurricane Lee, respectively. b) and f) display the error distribution of the OCN model at these two time steps, while c) and g) show the error distribution of the WW3 model at the same time steps. By comparison, it can be observed that the errors of OCN are generally smaller than those of WW3, especially near the hurricane center. d) presents the comparison curve of the ACC between the two models over the 11 time steps. The ACC of OCN is consistently higher than that of WW3, indicating that its predictions have a better correlation with the true values. (h) shows the comparison curve of the RMSE between the two models. The RMSE of OCN is consistently lower than that of WW3, suggesting that its prediction errors are smaller. Overall, OCN demonstrates higher prediction accuracy than WW3 in the simulation of Hurricane Lee.
During the 11-time-step (2.75 days) forecasting period, within the forecasting area (longitude 72°W to 30°W, latitude 0°N to 30°N), the MAE of the OCN model (0.17 m) was lower than that of the WW3 model (0.26 m). In terms of ACC, the minimum value for the OCN model (0.79) was slightly lower than that of the WW3 model (0.82). Regarding RMSE, the maximum value for the OCN model (0.22 m) was similar to that of the WW3 model (0.21 m). However, if we focus on the first 4 time steps (before the WW3 model updates its forecasting results), the minimum ACC value for the OCN model (0.97) was higher than that of the WW3 model (0.91), and the maximum RMSE value for the OCN model (0.13 m) was lower than that of the WW3 model (0.21 m). Based on these results, we can conclude that the short-term forecasting performance of the OCN model is comparable to that of traditional forecasting models.
3.3 Analysis of the Model’s Physical Properties
Given the good performance demonstrated above, we can conclude that the deep learning process is not merely a result of simple data statistics. The model has likely captured the intrinsic relationship between wind and waves in a fixed topographic environment. To further investigate this, we conducted the following two experiments:
3.3.1 normal conditions
In traditional models, after preparing accurate external forcing (i.e., the factors that influence the model’s behavior but are not directly simulated by the model, such as wind fields), the model’s simulation results will typically start from a relatively inaccurate state and reach a stable condition after a period of running. This process reflects the balance achieved by various physical processes represented in the model. To investigate whether our deep learning model exhibits similar characteristics, we used the ERA5 wind fields at 12:00 and 16:00 on February 11, 2023, as the initial conditions, setting the wave field to zero. We then used the ERA5 wind fields as input to simulate the wave field for 15 days (from February 12 to February 27). The results are shown in Figure 7.
Figure 7: Verification experiment of the OCN model’s wave generation process under zero initial conditions. a) and c) show the global distribution of the absolute error of the OCN model at time step 1 and time step 63, respectively. It can be seen from the figures that at time step 63, the error of the OCN model is small in most sea areas, indicating that the model can well simulate the wave generation process. b) and d) show the ACC and RMSE curves of the model prediction results with respect to the forecast time. As the forecast time increases, the ACC gradually improves and stabilizes around 0.85, while the RMSE gradually decreases and stabilizes around 0.55m, further confirming the excellent performance of the OCN model in simulating the wave generation process. Overall, this experiment demonstrates that the OCN model can reproduce the wave generation process very well with high physical rationality.
During the 63-time-step running process, the model gradually self-corrected, and the final results tended to stabilize. The ACC value increased from -0.08 to around 0.85, while the RMSE decreased from 2.36 m to around 0.55 m. These trends demonstrate that the deep learning model exhibits similar characteristics to traditional physical models in terms of self-correction and convergence to a stable state. Specifically, the ACC value, which represents the correlation between the simulated and observed wave fields, started from a negative value (-0.08) due to the initially zero wave field. As the model ran, the ACC value gradually increased, indicating that the simulated wave field became increasingly consistent with the observed wave field. Similarly, the RMSE, which measures the average magnitude of the errors between the simulated and observed wave fields, started from a relatively high value (2.36 m) and decreased over time, indicating that the model’s accuracy improved as it self-corrected. These results suggest that the deep learning model, like traditional physical models, can capture the intrinsic relationships between wind and waves and simulate the dynamic processes of wave generation, propagation, and dissipation. This similarity in behavior supports the conclusion that the deep learning model has learned the underlying physical mechanisms governing wind-wave interactions.
3.3.2 extreme situations
We also selected Mawar and Lee as comparison objects to simulate the entire process. First, for Mawar, we used May 18, 2023, as the starting point for the simulation. We employed the ERA5 wind fields at 12:00 and 16:00 on that day as the initial wind fields, setting the initial wave field to zero. The ERA5 wind fields were used as input, and the simulation was carried out until June 11. The results are shown in Figure 8.
Figure 8: Figure shows the simulation performance of the OCN model during a complete typhoon process. a) and d) show the 10-meter wind speed distribution in the early (time step 1) and late (time step 95) stages of the simulation, respectively, presenting the entire process of typhoon generation, development, and dissipation. b) and e) show the spatial distribution of the absolute error of the OCN model at time steps 1 and 95, respectively. Comparing the two figures, it can be found that as the simulation time progresses, the error of the OCN model decreases significantly. Even in high wind speed areas such as the typhoon center, the model can provide relatively accurate forecasts in the later stage, indicating that the model can well simulate the wave evolution during the typhoon process. c) and f) show the ACC and RMSE curves of the model prediction results with respect to the forecast time. It can be seen that as the forecast time increases, the prediction skill of the model continuously improves, with ACC stabilizing at a high value of around 0.96 and RMSE decreasing to a low value of around 0.26 meters, quantitatively confirming the excellent performance of the OCN model in typhoon wave simulation.
During the 95-time-step simulation process, the ACC for the entire typhoon region increased from -0.13 to around 0.96, while the RMSE decreased from 1.38 m to around 0.36 m. The increasing ACC indicates that the simulated wave field became increasingly consistent with the observed wave field, while the decreasing RMSE suggests that the model’s accuracy improved over time. From the last time step, it can be observed that the errors in most areas were within 1 m, with the main error areas located near the typhoon. These results demonstrate that even under the influence of extreme wind fields, the model exhibits a convergence process similar to that of physical models.
For Lee, we used September 4, 2023, as the starting point for the simulation. We employed the ERA5 wind fields at 12:00 and 16:00 on that day as the initial wind fields, setting the initial wave field to zero. The ERA5 wind fields were used as input, and the simulation was carried out until September 18. The results are shown in Figure 9.
Figure 9: Simulation performance of the OCN model during Hurricane Lee. a) and d) show the 10-meter wind speed distribution at simulation time steps 1 and 55, respectively. b) and e) show the spatial distribution of the absolute error of the OCN model at time steps 1 and 55, respectively, indicating that the model error decreases significantly as the simulation progresses. c) and f) show the ACC and RMSE curves of the model prediction results with respect to the forecast time, where ACC eventually stabilizes around 0.97 and RMSE decreases to around 0.26 meters, confirming the excellent performance of the OCN model in the wave simulation of Hurricane Lee.
During the 55-time-step simulation process, the ACC for the entire hurricane region increased from 0.52 to around 0.97, while the RMSE decreased from 1.78 m to around 0.26 m. Similar to the case of Mawar, these trends indicate that the model’s performance improved over time, with the simulated wave field becoming increasingly consistent with the observed wave field. From the last time step, it can be observed that, similar to the situation with Mawar, the errors in most areas were within 1 m, with the main error areas located near the hurricane. These results further confirm that even under the influence of extreme wind fields, the model exhibits a convergence process similar to that of physical models. From the above results, it can be concluded that, even in extreme situations, the model demonstrates a convergence process similar to that of physical models. We can consider that the model exhibits a physical model-like convergence effect even under extreme conditions.
4 Discussion
In the previous experiments, our wave model demonstrated that deep learning models can also exhibit physical characteristics very well. In fact, our model can be considered as a deep learning version of WW3. However, in recent years, deep learning meteorological models have encountered some troubling issues, especially during long-term runs, where the models exhibit significant energy dissipation states. This phenomenon piqued our interest because it stands in stark contrast to the energy conservation properties we observed in the wave model.
To better understand the energy dissipation problem in meteorological models, we selected FUXI [12], currently the best-performing deep learning meteorological model, and ECMWF Integrated Forecasting System (IFS), a traditional meteorological model commonly used for performance comparison, as our research objects. Considering that differences in spatial resolution between models may affect the total energy magnitude, we used energy density instead of total energy as the comparison metric. Specifically, we used wind energy density to describe the energy state of the meteorological models and wave energy density to describe the energy state of the wave model. By randomly selecting a day as the starting point, we compared the forecasting effects of FUXI and IFS after long-term runs. As shown in Figure 10, although FUXI’s ACC is significantly better than IFS, its wind energy density curve deviates far from the true values and exhibits a clear divergence state at the last time step. In contrast, our wave model maintained a consistent trend with the actual wave energy density changes throughout the forecasting process, and the forecasting results at the last time step still showed a state full of energy.The formulas for calculating wind energy density and wave energy density are detailed in the Appendix A.
The reason for this difference is that when constructing the wave model, we considered wind energy and previous wave energy as inputs, thereby maintaining energy balance within the model. In contrast, current deep learning meteorological models do not explicitly consider the issue of energy conservation. To further demonstrate the importance of energy balance for improving model performance, we attempted to construct a simple meteorological model, OceanCastNet-wind (OCN-wind). This model takes the 10-m wind field at the sea surface as the forecasting object and selects factors that have a significant impact on wind field energy changes as input variables, including the wind fields at the current and previous time steps, as well as solar radiation intensity at the current, previous, and future time steps. Specifically, the input-output relationship of the model can be represented by the following function:
$[u_{10}(t+1),v_{10}(t+1)]=g[R(t-1),u_{10}(t-1),v_{10}(t-1),R(t),u_{10}(t),v_{10}(t),R(t+1)],$ (2)
where $u_{10}$ and $v_{10}$ represent the zonal and meridional wind speeds at 10 m height, respectively, $R$ represents the incident solar radiation flux at the top of the atmosphere (TOA), and $g$ represents the model’s mapping function.
We used the same training data and methods as OCN to train OCN-wind and compared its long-term forecasting results with FourCastNet, which is also based on adaptive Fourier operators. As shown in Figure 11, at the 75th forecasting time step, FourCastNet’s forecasting results have already exhibited significant energy dissipation phenomena, with some areas even experiencing numerical instability issues. In contrast, the forecasting results of OCN-wind always maintain a reasonable energy level without any signs of collapse. This result proves that by introducing energy balance constraints into the model, the long-term forecasting performance of deep learning meteorological models can be effectively improved.
Figure 10: Comparison of wind speed prediction performance among different models. a) and d) show the 10-meter wind speed ACC curves of the ECMWF Integrated Forecasting System (IFS) and FUXI models, respectively, with black and gray lines representing u and v components. Although FUXI’s ACC consistently outperforms IFS, the actual prediction results c) and f) reveal that FUXI’s predictions exhibit rapid energy dissipation, while IFS’s predictions are closer to the actual situation. b) and e) present the wind energy density curves of the two models, with blue lines representing model predictions and red lines representing ERA5 reanalysis data. It can be observed that the wind energy density of IFS is highly consistent with ERA5, while FUXI shows significant deviations. This might be because current meteorological deep learning models are mostly autoregressive, lacking continuous energy input (such as solar radiation) during the continuous forecasting process, resulting in continuous energy dissipation. In contrast, the OCN model, due to the continuous wind (energy) input, can maintain a relatively high ACC g), a wave energy density curve close to the actual situation (h), and relatively less dissipation characteristics i) in the later stage of the forecast.
Figure 11: Comparison of energy evolution in continuous forecasting between different models. To verify the impact of continuous energy input on model forecasting performance, we modified the wave model OCN into the OceanCastNet-wind (OCN-wind) model, enabling it to receive TOA incident solar radiation as an additional input condition. Figures a) and b) show the forecasting results of the FourCastNet (FCN) model at time steps 1 and 75, respectively, while figures c) and d) show the forecasting results of the OCN-wind model at the same time steps. It can be observed that although the FCN model exhibits significant energy dissipation in the later stage of forecasting, the OCN-wind model can still maintain energy characteristics relatively well. This result confirms that continuous energy input plays an important role in improving the long-term forecasting performance of deep learning models, providing new ideas for subsequent model improvements.
5 Conclusion
In this study, we developed a deep learning-based wave forecasting model OceanCastNet (OCN) and compared its performance with a traditional physical model, WaveWatch III (WW3). Our results demonstrate that the OCN model can effectively learn the intrinsic relationships between wind and waves, capturing the complex physical processes involved in wave generation, propagation, and dissipation.
Through a series of experiments, we found that the OCN model exhibits superior performance compared to traditional models in both normal and extreme conditions. In the comparison with the WW3 model, OCN showed lower mean absolute errors, higher anomaly correlation coefficients, and lower root mean square errors, indicating its ability to provide more accurate wave forecasts. Furthermore, the OCN model demonstrated a convergence process similar to that of physical models, even under the influence of extreme wind fields, such as those associated with typhoons and hurricanes.
Our analysis of the energy density in the model results revealed that the deep learning process not only learns the statistical relationships between variables but also captures the underlying physical mechanisms. By comparing the wind energy density and wave energy density curves of different models, we found that the OCN model maintains a better balance between energy input and dissipation, closely resembling the actual observation results. In contrast, some autoregressive models, such as FUXI and FourCastNet, exhibited significant energy dissipation over time, highlighting the importance of incorporating energy input in long-term forecasting.
Moreover, our study sheds light on the potential limitations of current autoregressive models in meteorological and oceanographic applications. These models often rely solely on the previous time step’s forecast results, neglecting other crucial inputs, such as solar radiation, which drives atmospheric and oceanic circulations. By incorporating energy input terms, as demonstrated in the OCN-wind model, we showed that deep learning models can maintain more realistic wind field characteristics over extended simulation periods.
In conclusion, our findings suggest that deep learning models, such as OCN, have the potential to revolutionize wave forecasting and other geophysical applications. By learning the intrinsic physical processes and incorporating essential energy input terms, these models can provide accurate and reliable long-term predictions, outperforming traditional physical models and existing deep learning approaches. Future research should focus on further improving the representation of physical processes in deep learning models and exploring their applicability to a wider range of geophysical phenomena.
Acknowledgments
We would like to express our gratitude to the European Centre for Medium-Range Weather Forecasts (ECMWF) for providing the ERA5 reanalysis data used for model training. The WaveWatch III (WW3) model data used for comparison were obtained from the Pacific Islands Ocean Observing System (PacIOOS), and the wind field input data were provided by Dr. Kwok Fai Cheung and his lab at the Department of Ocean and Resources Engineering, University of Hawaii at Mānoa. The FCN forecasting results were generated using the code provided by the ECMWF Lab, while the IFS results were obtained from the TIGGE archive. The FUXI forecasting results were produced using the code kindly provided by the authors. We also acknowledge the National Oceanic and Atmospheric Administration (NOAA) National Centers for Environmental Prediction (NCEP) for providing the Global Forecast System (GFS) data.
We would like to thank the Dawning Information Industry Co., Ltd. for providing the computational resources used in this study.
We extend our special thanks to the authors of FourCastNet for their inspiring work. The architecture of our OceanCastNet model is built upon the framework of FourCastNet, and we are grateful for their innovative contributions to the field of deep learning-based weather forecasting.
Appendix A Appendix: Formulas for Calculating ACC, RMSE, Wind Power Density, and Wave Power Density
We employed the latitude-weighted anomaly correlation coefficient (ACC) and root mean square error (RMSE) to evaluate the model performance. The definitions of these metrics are derived from [13], but we adopted the slightly modified versions from [4]. The latitude-weighted ACC is defined as follows:
$ACC(v,l)=\frac{\sum_{m,n}L(m)\tilde{X}{pred}{(l)}[v,m,n]\tilde{X}{true}{(l)}[v,m,n]}{\sqrt{\sum_{m,n}L(m)(\tilde{X}{pred}{(l)}[v,m,n])^{2}\sum_{m,n}L(m)(\tilde{X}_{true}{(l)}[v,m,n])^{2}}},$ (3)
where $\tilde{X}{pred/true}{(l)}$ represents the long-term-mean-subtracted value of predicted (/true) variable $v$ at the location denoted by the grid coordinates $(m,n)$ at the forecast time step $l$. The long-term mean of a variable is simply the mean value of that variable over a large number of historical samples in the training dataset. The long-term mean-subtracted variables $\tilde{X}{pred/true}$ represent the anomalies of those variables that are not captured by the long-term mean values. $L(m)$ is the latitude weighting factor at the coordinate $m$, defined as follows:
$L(j)=\frac{\cos(lat(m))}{\frac{1}{N_{lat}}\sum_{j}^{N_{lat}}\cos(lat(m))}.$ (4)
The latitude-weighted RMSE for a forecast variable $v$ at forecast time step $l$ is defined by the following equation, using the same latitude weighting factor as ACC:
$RMSE(v,l)=\sqrt{\frac{1}{NM}\sum_{m=1}^{M}\sum_{n=1}^{N}L(m)(X_{pred}{(l)}[v,m,n]-X_{true}{(l)}[v,m,n])^{2}},$ (5)
where $X_{pred/true}^{(l)}[v,m,n]$ represents the value of predicted (/true) variable $v$ at the location denoted by the grid coordinates $(m,n)$ at the forecast time step $l$.
We use the following formula to calculate the wind energy density in the operation results:
$E_{\text{wind}}(l)=\frac{1}{NM}\sum_{m=1}^{M}\sum_{n=1}^{N}\frac{1}{2}\rho_{\text{air}}\left(u_{10}^{2}(l)[m,n]+v_{10}^{2}(l)[m,n]\right)^{3/2}$ (6)
where $u_{10}$ and $v_{10}$ represent the wind speed at 10 m above the sea surface in the model operation results, and $\rho_{\text{air}}$ is the air density, taken as 1.225 kg/m³ here.
We use the following formula to calculate the wave energy density:
$E_{\text{wave}}(l)=\frac{1}{NM}\sum_{m=1}^{M}\sum_{n=1}^{N}\frac{1}{16}\rho_{\text{sea}}gH^{2}(l)[m,n]$ (7)
where $\rho_{\text{sea}}$ is the seawater density, which we take as 1,025 kg/m³, $g$ is the gravitational acceleration, taken as 9.8 m/s² here, and $H$ is the significant wave height.